OpenStudy (anonymous):

Annie wrote the equation shown below. 3(3y) = (3 • 3) y Which property did Annie apply? associative property of multiplication commutative property of multiplication multiplication property of equality distributive property of multiplication

OpenStudy (anonymous):

Associativity A⋆(B⋆C)=(A⋆B)⋆C Commutativity A⋆B=B⋆A Distributitivity A⋅(B⋆C)=(A⋅B)⋆(A⋅B)

OpenStudy (anonymous):

its not multi... eliminate that

OpenStudy (anonymous):

Associative property of multiplication says that : \[a \cdot (b \cdot c) = (a \cdot b) \cdot c\]

OpenStudy (anonymous):

so its associative you have proof

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Ha ha ha ha..

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

want to help with one more?

OpenStudy (anonymous):

Frank writes the statement shown below. “The sum of 4 and 6 is the same as 2 times the sum of 2 and 3.” Which property of numbers did Frank apply to write the statement?

OpenStudy (anonymous):

addition property of equality associative property of addition commutative property of addition distributive property of addition

OpenStudy (anonymous):

2(2+3)=4+6

OpenStudy (anonymous):

\[2(2 + 3) = 2 \times 2 + 2 \times 3 = 4 + 6\]

OpenStudy (anonymous):

Associativity A⋆(B⋆C)=(A⋆B)⋆C Commutativity A⋆B=B⋆A Distributitivity A⋅(B⋆C)=(A⋅B)⋆(A⋅B)

OpenStudy (anonymous):

so associative?